hi, any help for this
Admission tickets to a theater were P60 for adults and P25 for children. Receipts for the day showed that 280 persons attended and P14,000 were collected. How many children attended that day?

Thanks
JM

Answer by scott8148(6628)

a + c = 280 ___ a = 280 - c

60a + 25c = 14000

substituting ___ 60(280 - c) + 25c = 14000 ___ 16800 - 60c + 25c = 14000

2800 = 35c